OpenStudy (anonymous):

Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

Y = 9 - 2x

OpenStudy (anonymous):

plug in y into the 3y - 2x = 11 equasion 3( 9 - 2x) - 2x = 11

OpenStudy (anonymous):

3*9 = 27 3*(-2x) = -6x 27 - 6x - 2x

OpenStudy (anonymous):

27 - 8x = 11 -27 -27 -8x = - 16 x = (-16)/(-8)

OpenStudy (anonymous):

x = 2 plug x in 3y -2x = 11 to solve for y 3y - 2(2) = 11

OpenStudy (anonymous):

3y - 4 = 11 +4 +4 3y = 15 y = 15/ 3 y = 5

OpenStudy (anonymous):

so x = 2 and y = 5 :)